OpenStudy (anonymous):

a flagpose is next to a building. the flagpole is 30 feet tall and casta a sahadow of 120 inches. the building casts a shadow of 648 inches long. how tall is the building in feet?

OpenStudy (anonymous):

flag pole is 30 feet tall, it's shadow is 120/12 = 10 feet. the ratio is 3:1 the shadow of the building is 648/12 = 54 feet 54*3=162 feet tall.

OpenStudy (anonymous):

thank you i got it right